# Doc2SAR: A Synergistic Framework for High-Fidelity Extraction of Structure-Activity Relationships from Scientific Documents

### v1

@@ 4.3 Key Components Analysis: The impact of SFT
We evaluate this task on our custom Coreference Val Set using Recall (%) as the metric. Our SFT-enhanced model (Ours) achieves an overall recall of 76.40%, significantly outperforming both GPT-4o (53.40%) and its base model, Qwen2.5-VL-7B-Ins. (36.60%). The improvement is most pronounced in the unstructured Mol-outside-Table scenario (70.45% vs. 38.81% for GPT-4o). In the more structured Mol-inside-Table setting, our model also remains highly competitive (95.76%), demonstrating its robustness across different layouts. This confirms that domain-specific SFT is crucial for handling the varied contexts in which molecules appear.
Table 2: Performance comparison on the Molecular Coreference Recognition task (Recall %).
@@
Overall Mol-inside-Table Mol-outside-Table
@@
We evaluate this task on the standard PubTabNet Val Set using the TEDS score (%) [29]. The results again highlight the power of SFT. Our model achieves a superior overall TEDS score of 58.29%, compared to 48.05% for GPT-4o and 43.51% for the base model. Its advantage is largest on tables with Complex Layouts (54.44% vs. 43.31% for GPT-4o), which involve challenging merged cells. Even on Standard Grids, our model maintains a clear lead (61.72% vs. 52.57% for GPT-4o), indicating an all-around improvement. This enhanced structural understanding is a key factor in the high performance of our Doc2SAR framework.
@@
Overall Standard Grids Complex Layouts
@@
5 Case Study
To provide qualitative evidence of our framework’s real-world efficacy, this section presents two representative case studies from diverse scientific documents. These cases demonstrate Doc2SAR’s ability to accurately process varied layouts and correctly associate molecular structures with their activity data, even under challenging conditions.
Figure 5: Case study in literature 10.1021_acsmedchemlett.3c00392. Note that the red and blue boxes (masks) in the case are only for auxiliary visualization display and are not included in the document. Doc2SAR correctly performs intra-page association, linking a molecular structure to its corresponding activity data table on the same page.
Figure 6: Case study in patent CN105793260B. Note that the red and blue boxes (masks) in the case are only for auxiliary visualization display and are not included in the document. Doc2SAR demonstrates its robustness by performing a long-range, cross-page association, linking a molecule on Page 34 with its activity data in a table on Page 70.
The first case, illustrated in Figure 5, involves a typical scientific literature article where the molecular structure and its corresponding activity data are located on the same page. Our framework successfully performs this intra-page association, correctly identifying the molecule, locating its entry in the activity table, and extracting the complete data row.
The second case, shown in Figure 6, highlights a more challenging scenario from a Chinese patent. The molecular structure is found on Page 34, while its activity data is in a table 36 pages later on Page 70. This large, cross-page span poses a significant challenge for monolithic end-to-end models. In contrast, our modular Doc2SAR framework excels, robustly linking the disparate information via its coreference-based merging strategy. This case also showcases the framework’s versatility in handling non-English documents.
@@ 6 Related Work
Early rule-based systems could recognise chemical names or parse fixed reaction templates, but generalised poorly to diverse layouts. Recent efforts have shifted toward neural, modality-specific extractors. For patents, the ChEMU2022 shared task released a benchmark for chemical NER and event extraction with layout cues [30]. Transformer OCSR models—DECIMER.ai with image–to–sequence decoding [31], MolScribe’s image-to-graph generation [32], and MolGrapher’s detection-plus-graph pipeline [33]—now achieve >90% top-1 exact match on curated sets. PatCID further demonstrated large-scale extraction of 80 M structure images from patents by combining segmentation with OCSR [34]. However, these systems operate on a single input type (text or image) and therefore cannot recover cross-modal SARs.
@@
Large language models are increasingly adapted for chemistry and materials science. ChemCrow [35] couples GPT-4 with cheminformatics tools to plan multi-step syntheses; Vangala et al. [36] fine-tuned GPT-J to mine patent reactions, boosting recall by 26 % over grammar-based baselines; FuncFetch [37] uses GPT-3.5 to extract >20 k enzyme–substrate pairs with high precision; and SynAsk [38] combines a knowledge base with a GPT-derived model for interactive retrosynthesis and reaction-condition suggestions. Although these LLMs excel at text-centric tasks, evaluations show degraded accuracy once molecular diagrams or numeric tables must be aligned with textual mentions [39]. Doc2SAR addresses this limitation by integrating fine-tuned multimodal LLMs with domain-specific tools to link text, figures and tables in documents.
@@
Vision–language document models fuse layout and visual cues: LayoutLMv3 [40] unifies text–image masking for pre-training, UReader [41] adopts an OCR-free encoder for long documents, and DocOwl 1.5 [42] refines multimodal alignment for page-level QA. Benchmarks such as UniMMQA [43] and SPIQA [44] probe reasoning across text, tables and figures, yet contain few chemical diagrams and no ground-truth SAR links. Therefore, DocSAR-200 contributes the first dataset in which molecular structures and activity tables are jointly annotated, and Doc2SAR provides a framework that demonstrably outperforms general VLMs on this document-level SARs extraction.
@@ 7 Conclusion
In this work, we addressed the critical challenge of automated SAR extraction from complex scientific documents. We introduced DocSAR-200, a comprehensive new benchmark comprising 200 diverse documents, specifically designed to drive progress in this domain. Concurrently, we proposed Doc2SAR, a novel synergistic framework integrating domain-specific tools with advanced MLLMs which enhanced by SFT. Our experiments demonstrate that the SFT-enhanced framework achieves a state-of-the-art Table Recall of 80.78%, significantly outperforming the strongest end-to-end baseline by a remarkable 51.48% over GPT4o. This result highlights the superiority of our modular design by demonstrating its ability to excel at fine-grained sub-tasks where general models falter, while also affirming the crucial role of SFT in adapting models for precise, domain-specific tasks. By providing robust tools and a standardized framework, our work helps accelerate the transition from manual curation to automated knowledge discovery in the sciences.
While Doc2SAR sets a new standard, we acknowledge current limitations, such as a primary focus on English documents and the need for curated data for SFT, which define clear avenues for future work:
  • Benchmark Expansion: We plan to expand DocSAR-200 with more languages (e.g., Chinese, German), diverse scientific domains, and more challenging annotation types.
  • Methodological Advancements: We will investigate hybrid synergistic models and explore more data-efficient fine-tuning techniques to reduce dependency on supervised data.
  • Broader Applications: We aim to adapt our framework to other critical scientific tasks, such as extracting experimental protocols, to foster a more comprehensive AI-driven ecosystem.
We believe that DocSAR-200 will serve as a valuable resource for the community, spurring the development of more sophisticated and accurate models for scientific document understanding. Doc2SAR provides a strong baseline and a flexible architecture that can be built upon for future research in this rapidly evolving field.
Appendix A Prompts
@@
Prompt for extracting activity data from table image System Message: You are an expert in the field of pharmaceutical chemistry, and your task is to extract activity data from an image in tabular HTML format.
User Message:
@@
Use the following guidelines:
1. Determine if the table has a header row. If the first row is a header, represent its cells with `<th> `elements. Otherwise, represent all rows (including the first row) with `<td> `elements.
2. Include a <table> element containing all rows (`<tr> `) and cells (`<th> `for headers, `<td> `for data).
3. Retain the exact content of each cell, including special characters, numerical values, text and even empty cell. 4. If a cell contains molecular structure images, replace it with the token "[mol]".
5. If the table contains merged cells, use appropriate attributes like `rowspan `or `colspan `to reflect the merging.
6. Pay attention to the following content: (1) Attribute: [’EC50’, ’IC50’, ’Ki’, ’Kd’, ’pKd’, ’TD50’, ’Ti’, ’TC50’]
(2) Unit: [’uM’, ’nM’, ’%’, ’kcal/mol’] If none of these Attribute keywords are found, it is not activity table. Your entire output should be the exact string: <table>None</table>
Example HTML structure:
@@
<table>
<tr>
<th>xxx</th>
<th>xxx</th>
<th>xxx</th>
</tr>
<tr>
<td>xxx</td>
<td>x</td>
<td>x</td>
</tr>
<tr>
<td>xxx</td>
<td>x</td>
<td>xx</td>
</tr>
</table>
[Image of a table may contains activity data]
Prompt for recognizing coreference from molecular structure image (In Case of Outside the table) System Message: You are an expert in the field of pharmaceutical chemistry, and your task is to recognize the coreference from a molecular structure image in json format.
User Message:
@@
Your task is to **extract only the molecular coreference** associated with the molecule. The coreference is typically a unique identifier (e.g., numeric, alphabetic, or alphanumeric code) found in the image.
Important Instructions:
1. **Focus on coreference priority**:
- **First, check the left side of the molecule** for a valid coreference.
- **If the left side is empty or invalid, check the right side.**
- Do **not** check inside any table.
2. **Ignore the molecular structure itself**:
- Do not output SMILES, InChI, or any chemical structure representation.
- Focus only on extracting the **molecular coreference**.
3. Strictly Follow The JSON Output Format
Example Json structure:
@@ 3. Strictly Follow The JSON Output Format
{{"compound_id": "<Molecular coreference or [None]>"}}
[Image of a molecular structure contains corresponding coreference]
Prompt for recognizing coreference from molecular structure image (In Case of Inside the table) System Message: You are an expert in the field of pharmaceutical chemistry, and your task is to recognize the coreference from a molecular structure image in json format.
User Message:
@@
Your task is to **identify only the molecular coreference** in the image, which is typically a unique identifier (e.g., a numeric, alphabetic, or abbreviations) associated with the molecule in the image.
Important Instructions:
1. **Priority of Coreference Extraction**:
- The molecular coreference **is usually inside the same table cell** as the molecular image.
- If the table cell is empty or ambiguous, check nearby table cells.
- Do **not** look outside the table.
2. **Ignore the molecular structure itself**:
- Do not output SMILES, InChI, or any chemical structure representation.
- Focus only on extracting the **molecular coreference**.
3. Strictly Follow The JSON Output Format
Example Json structure:
@@
{{"compound_id": "<Molecular Coreference or [None]>"}}
[Image of a molecular structure contains corresponding coreference]
@@ Appendix B Annotation Details and Quality Control
The DocSAR-200 annotation process was designed to ensure high fidelity and reproducibility. Each document was processed through a two-stage pipeline: first, GPT-4o’s multimodal capabilities [17] were employed to extract preliminary tabular information (e.g., compound names, activity values) into structured formats. In parallel, molecular structure images were located using coreference links or positional heuristics.
Subsequently, trained annotators reviewed and corrected all entries, ensuring accurate alignment between molecules, labels, and extracted tables. Molecular images were parsed into SMILES representations via open-source OCSR tools, such as MolParse [20], followed by human validation.
@@
Validation protocols. All SMILES were validated using RDKit [45] to ensure chemical correctness and canonicalization. Heuristic filters removed malformed structures and disconnected fragments. Coordinates and page indices were manually verified to prevent drift or misalignment.
@@ Appendix C SFT Dataset Construction Details
SFT dataset is meticulously constructed from a combination of real-world and synthetic molecular images, designed to enhance the model’s capabilities in recognizing molecular coreferences. To create a unified training corpus, all samples are processed into one of two standardized layouts: an in-table layout, which captures the entire table row containing the target molecule, and an out-of-table layout, where the molecule image is cropped with fixed padding to preserve essential surrounding context.
The real data samples originate from a diverse collection of patents and scientific PDFs, reflecting authentic, "in-the-wild" scenarios. Our semi-automated annotation pipeline begins with UniFinder to detect the bounding boxes of molecules, followed by PyMuPDF [46] to crop the corresponding images based on their layout type. An initial annotation pass is performed by Qwen2.5-VL-7b-Instruct [24] using domain-specific prompts (see Appendix A), with all generated labels subsequently undergoing a rigorous two-person verification process to ensure accuracy. To further augment the dataset, especially for covering edge cases and creating challenging negative samples, we generate synthetic data. These samples are created by rendering molecular structures from PubChem [26] into both layouts. Their labels are randomly generated based on real-data formats, and for out-of-table samples, label positions are programmatically assigned using OpenCV [47] to approximate empirical distributions (e.g., bottom: 50%, top: 30%).
The final consolidated SFT dataset comprises a total of 9,790 samples. This includes 4,594 real-world samples, of which 1,747 are in the in-table layout and 2,847 are out-of-table. The dataset is further enriched by 5,196 synthetic samples, composed of 510 in-table instances and a large number of out-of-table hard cases (4,686), specifically designed to enhance the model’s robustness against complex visual contexts. Figure 7 provides a visual summary of these distributions.
Figure 7: Statistical composition of the SFT dataset for the Mol ID agent. The chart displays absolute sample counts for: (i) In-table and out-of-table components within Real Data, Synthetic Data, and the Overall Dataset. (ii) A direct comparison of total Real Data versus total Synthetic Data under ’Source Totals’.

### v2

@@ 4.3 Components Analysis: The impact of SFT
We evaluate this task on our custom Coreference Recognition Val Set using Recall (%) as the metric. We categorize test scenarios into Simple cases, where molecules appear within structured table environments with clear spatial relationships, and Hard cases, where molecules are embedded in unstructured contexts such as figure captions, chemical schemes, or isolated paragraph text. Our SFT-enhanced model achieves an overall recall of 76.40%, substantially outperforming both GPT-4o (53.40%) and the base Qwen2.5-VL-7B-Instruct (36.60%). The improvement is particularly striking for Hard scenarios, where our model achieves 70.45% compared to GPT-4o’s 38.81%, an 82% relative improvement. Even for Simple cases, our model maintains excellent performance at 95.76%, demonstrating robust capabilities across diverse molecular presentation contexts.
Table 2: Performance comparison on the Molecular Coreference Recognition task, measured by Recall (%).
@@
Overall Simple Hard
@@
We evaluate this task on the standard PubTabNet Val Set using the TEDS score (%) metric. Tables are categorized as Simple (regular grid layouts with uniform cell structures and clear row-column organization) and Hard (complex layouts featuring merged cells, multi-level headers, nested structures, or irregular formatting). Our model achieves a superior overall TEDS score of 58.29%, outperforming GPT-4o (48.05%) and the base model (43.51%). The performance gain is most pronounced for Hard tables, where our model scores 54.44% versus GPT-4o’s 43.31%, a 26% relative improvement. This enhanced structural understanding of complex table layouts is crucial for accurately processing the diverse tabular formats found in scientific literature.
5 Case Study
To provide qualitative evidence of our framework’s real-world efficacy, we present two case studies showcasing intra-page vs. cross-page association and monolingual vs. multilingual document processing. These cases demonstrate Doc2SAR’s ability to handle diverse layouts and correctly associate molecular structures with activity data.
@@ 5 Case Study
Overall Simple Hard
@@
Qwen2.5-VL-7B-Instruct
The first case (Figure 5) involves a biomedical research paper where the molecular structure and activity data are co-located on the same page. Our framework successfully performs this intra-page association by: (1) extracting the molecular structure via OCSR, (2) identifying the compound’s textual identifier through coreference recognition, and (3) locating the matching table row. The system correctly extracts the IC50 value (2.3 nM), while GPT-4o failed due to OCSR errors.
The second case (Figure 6) presents a challenging scenario from a Chinese patent, where the molecular structure appears on Page 34 while activity data resides in a table 36 pages later. This cross-page span, combined with patent formatting and non-English text, overwhelms end-to-end approaches. Our framework demonstrates robust performance by accurately recognizing the molecular structure, extracting the compound identifier via multilingual coreference recognition, and successfully linking across the 36-page gap using fuzzy matching (similarity score: 0.95). This demonstrates scalability and language-agnostic capabilities.
@@ 6 Related Work
Early rule‐based systems could recognise chemical names or parse fixed reaction templates, but generalised poorly to diverse layouts. Recent efforts have shifted toward neural, modality‐specific extractors. For patents, the ChEMU2022 shared task released a benchmark for chemical NER and event extraction with layout cues (Li et al. 2022a). Transformer OCSR models include DECIMER.ai with image–to–sequence decoding (Rajan et al. 2024), MolScribe’s image-to-graph generation (Qian et al. 2023), and MolGrapher’s detection-plus-graph pipeline (Morin, Danelljan et al. 2023). These models now achieve more than 90% top-1 exact match on curated sets. PatCID further demonstrated large-scale extraction of 80 M structure images from patents by combining segmentation with OCSR (Morin et al. 2024). However, these systems operate on a single input type (text or image) and therefore cannot recover cross-modal SARs.
Figure 5: Case study in ACS Medicinal Chemistry Letters 10.1021_acsmedchemlett.3c00392. Note that the red and blue boxes (masks) in the case are only for auxiliary visualization display and are not included in the document. Doc2SAR correctly performs intra-page association, linking a molecular structure to its activity table on the same page.
@@
Large language models are increasingly adapted for chemistry and materials science. ChemCrow (Bran et al. 2023) couples GPT-4 with cheminformatics tools to plan multi-step syntheses; Vangala et al. (Vangala, Krishnan et al. 2024) fine-tuned GPT-J to mine patent reactions, boosting recall by 26 % over grammar-based baselines; FuncFetch (Smith et al. 2024) uses GPT-3.5 to extract >> 20 k enzyme–substrate pairs with high precision; and SynAsk (Zhang, Lin et al. 2025) combines a knowledge base with a GPT-derived model for interactive retrosynthesis and reaction-condition suggestions. Although these LLMs excel at text-centric tasks, evaluations show degraded accuracy once molecular diagrams or numeric tables must be aligned with textual mentions (Dagdelen, Dunn et al. 2024). Doc2SAR addresses this limitation by integrating fine-tuned multimodal LLMs with domain-specific tools to link text, figures and tables in documents.
@@
Vision language document models fuse layout and visual cues: LayoutLMv3 (Huang et al. 2022) unifies text–image masking for pre-training, UReader (Ye et al. 2023) adopts an OCR-free encoder for long documents, and DocOwl 1.5 (Hu et al. 2024) refines multimodal alignment for page-level QA. Benchmarks such as UniMMQA (Luo, Shen, and Deng 2023) and SPIQA (Pramanick, Chellappa, and Venugopalan 2024) probe reasoning across text, tables and figures, yet contain few chemical diagrams and no ground-truth SAR links. Therefore, DocSAR-200 contributes the first dataset in which molecular structures and activity tables are jointly annotated, and Doc2SAR provides a framework that demonstrably outperforms general VLMs on this document-level SARs extraction.
Figure 6: Case study in C.N. patent CN105793260B. Note that the red and blue boxes (masks) in the case are only for auxiliary visualization display and are not included in the document. Doc2SAR demonstrates its robustness by performing a long-range, cross-page association, linking a molecule on Page 34 with its activity data on Page 70.
@@ 7 Conclusion
In this work, we introduce DocSAR-200, a benchmark of 200 scientific documents, and Doc2SAR, a synergistic framework integrating domain-specific tools with SFT-enhanced MLLMs for automated SAR extraction. Doc2SAR achieves state-of-the-art Table Recall of 80.78%, a 175% relative improvement over the strongest end-to-end baseline, demonstrating that specialized architectures outperform monolithic approaches.
Our synergistic framework demonstrates that strategic integration of specialized components yields superior performance over large general-purpose models. By enabling each component to excel at its designated task while maintaining seamless cross-modal coordination, this modular paradigm fundamentally addresses the limitations inherent in monolithic approaches to scientific document understanding.
Future work will explore extending multilingual capabilities, developing data-efficient methods, and adapting the framework to other tasks. We anticipate DocSAR-200 will catalyze research in scientific document understanding.
Appendix A Prompts Design
@@
Prompt for extracting activity data from table image System Message: You are an expert in the field of pharmaceutical chemistry, and your task is to extract activity data from an image in tabular HTML format. User Message:
@@
Use the following guidelines: 1. Determine if the table has a header row. If the first row is a header, represent its cells with `<th>` elements. Otherwise, represent all rows (including the first row) with `<td>` elements. 2. Include a <table> element containing all rows (`<tr>`) and cells (`<th>` for headers, `<td>` for data). 3. Retain the exact content of each cell, including special characters, numerical values, text and even empty cell. 4. If a cell contains molecular structure images, replace it with the token ”[mol]”. 5. If the table contains merged cells, use appropriate attributes like `rowspan ` or `colspan ` to reflect the merging. 6. Pay attention to the following content: (1) Attribute: [’EC50’, ’IC50’, ’Ki’, ’Kd’, ’pKd’, ’TD50’, ’Ti’, ’TC50’] (2) Unit: [’uM’, ’nM’, ’%’, ’kcal/mol’] If none of these Attribute keywords are found, it is not an activity table. Your entire output should be the exact string: <table>None</table> Example HTML structure:
@@
<table> <tr> <th>xxx</th> <th>xxx</th> <th>xxx</th> </tr> <tr> <td>xxx</td> <td>x</td> <td>x</td> </tr> <tr> <td>xxx</td> <td>x</td> <td>xx</td> </tr> </table> ``` [Image of a table may contain activity data]
Prompt for recognizing coreference from molecular structure image (In Case of Outside the table) System Message: You are an expert in the field of pharmaceutical chemistry, and your task is to recognize the coreference from a molecular structure image in JSON format. User Message:
@@
Your task is to **extract only the molecular coreference** associated with the molecule. The coreference is typically a unique identifier (e.g., numeric, alphabetic, or alphanumeric code) found in the image. Important Instructions: 1. **Focus on coreference priority**: - **First, check the left side of the molecule** for a valid coreference. - **If the left side is empty or invalid, check the right side.** - Do **not** check inside any table. 2. **Ignore the molecular structure itself**: - Do not output SMILES, InChI, or any chemical structure representation. - Focus only on extracting the **molecular coreference**. 3. Strictly Follow The JSON Output Format Example JSON structure:
@@
{{”compound_id”: ”<Molecular coreference or [None]>”}} ``` [Image of a molecular structure containing corresponding coreference]
Prompt for recognizing coreference from molecular structure image (In Case of Inside the table) System Message: You are an expert in the field of pharmaceutical chemistry, and your task is to recognize the coreference from a molecular structure image in JSON format. User Message:
@@
Your task is to **identify only the molecular coreference** in the image, which is typically a unique identifier (e.g., a numeric, alphabetic, or abbreviation) associated with the molecule in the image. Important Instructions: 1. **Priority of Coreference Extraction**: - The molecular coreference **is usually inside the same table cell** as the molecular image. - If the table cell is empty or ambiguous, check nearby table cells. - Do **not** look outside the table. 2. **Ignore the molecular structure itself**: - Do not output SMILES, InChI, or any chemical structure representation. - Focus only on extracting the **molecular coreference**. 3. Strictly Follow The JSON Output Format Example JSON structure:
@@
{{”compound_id”: ”<Molecular Coreference or [None]>”}} ``` [Image of a molecular structure containing corresponding coreference]
Figure 7: Composition of the SFT dataset. Left: breakdown of real and synthetic samples by layout type (in-table vs. out-of-table). Center: train (500 samples) and test (5,873 samples) splits, showing contributions from real and synthetic sources. Right: image size distribution categorized into Small, Medium, and Large, with respective sample counts.
@@ Appendix B Annotation Details and Quality Control
The DocSAR-200 annotation process was designed to ensure high fidelity and reproducibility. Each document was processed through a two-stage pipeline: first, GPT-4o’s multimodal capabilities were employed to extract preliminary tabular information (e.g., compound names, activity values) into structured formats. In parallel, molecular structure images were located using coreference links or positional heuristics.
Subsequently, trained annotators reviewed and corrected all entries, ensuring accurate alignment between molecules, labels, and extracted tables. Molecular images were parsed into SMILES representations via open-source OCSR tools, such as MolParse , followed by human validation.
@@
Validation protocols. All SMILES were validated using RDKit to ensure chemical correctness and canonicalization. Heuristic filters removed malformed structures and disconnected fragments. Coordinates and page indices were manually verified to prevent drift or misalignment.
@@ Appendix C SFT Dataset Construction Details
The SFT dataset is meticulously constructed from a combination of real-world and synthetic molecular images, designed to enhance the model’s capabilities in recognizing molecular coreferences. To create a unified training corpus, all samples are processed into one of two standardized layouts: an in-table layout, which captures the entire table row containing the target molecule, and an out-of-table layout, where the molecule image is cropped with fixed padding to preserve essential surrounding context.
The real data samples originate from a diverse collection of patents and scientific PDFs, reflecting authentic, “in-the-wild” scenarios. Our semi-automated annotation pipeline begins with UniFinder to detect the bounding boxes of molecules, followed by PyMuPDF to crop the corresponding images based on their layout type. An initial annotation pass is performed by Qwen2.5-VL-7b-Instruct using domain-specific prompts (see Appendix A), with all generated labels subsequently undergoing a rigorous two-person verification process to ensure accuracy. To further augment the dataset, especially for covering edge cases and creating challenging negative samples, we generate synthetic data. These samples are created by rendering molecular structures from PubChem into both layouts. Their labels are randomly generated based on real-data formats, and for out-of-table samples, label positions are programmatically assigned using OpenCV to approximate empirical distributions (e.g., bottom: 50%, top: 30%).
The final consolidated SFT dataset comprises a total of 6,373 samples. This includes 5,552 real-world samples (3,685 in in-table layout and 1,867 out-of-table) and 821 synthetic samples (532 in-table and 289 out-of-table). The dataset is further divided into 437 training samples and 5,873 testing samples.
Additionally, we analyze image size distributions to better understand the dataset characteristics. As shown in Figure 7, image sizes are categorized into three groups: Small (0–598,590 px2, 3,187 samples), Medium (598,590–927,368 px2, 1,593 samples), and Large (¿927,368 px2, 1,593 samples). The image sizes range from a minimum of 191,205 px2 to a maximum of 2,098,539 px2, with a mean of 668,888.55 px2. Figure 7 provides a comprehensive visualization of the dataset statistics, including real/synthetic distributions, train/test splits, and image size categories.